OpenStudy (anonymous):

What is the inverse of h? h(x) = 6x + 1

OpenStudy (anonymous):

\(h\) says a) multiply by 6 b) add 1 inverse says a) subtract 1 b) divide by 6 \[h^{-1}(x)=\frac{x-1}{6}\]

OpenStudy (anonymous):

like if you were solving \[6x+1=17\] you would a) subtract 1 b) divide by 6 \[6x+1=17\] \[x=\frac{17-1}{6}\]

OpenStudy (anonymous):

umm this isnt one of the multiple choice answers